OpenStudy (anonymous):

Can someone please help solve the system of equations algebraically y=x^2+2x y=3x+20

OpenStudy (janu16):

substitution?

OpenStudy (anonymous):

would I do 2x+x^2=20+3x ? to find x?

OpenStudy (phi):

2x+x^2=20+3x yes. now write that in standard form. Notice you should get a "quadratic" and standard form is a x^2 + bx + c = 0 where a,b, c are "coefficients" (numbers)